OpenStudy (anonymous):

In a 30° - 60° - 90° triangle, the longer leg is which of the following? (Points : 1) a) 3 times the length of the shorter leg b) 2 times the length of the hypotenuse c) 2 times the length of the shorter leg d) 3 times the length of the hypotenuse

OpenStudy (anonymous):

A 30o - 60o - 90o triangle is formed by splitting an equilateral triangle in two. If the equilateral triangle has side lengths of 2, that means the bottom triangle leg now was a length of 1, with a hypoteneuse of 2. By using the Pythagorean theorem, we find that the remaining side is root 2. Thus, the answer is root 3 times the length of the shorter leg.

OpenStudy (anonymous):

does that help? @allibricker12

OpenStudy (anonymous):

thank you so much! @kittenlover731

OpenStudy (anonymous):

what about this? @kittenlover731 In a 30° - 60° - 90° right triangle, the hypotenuse has a length of 8. What is the length of the longer leg? a) 4 b) 8 c) 4\[\sqrt{3}\] d) 4\[\sqrt{2}\]

OpenStudy (anonymous):

The length of the two legs are, shorter leg = 8*sin30 and longer leg = 8sin60 sin60 = √3/2 Longer leg = 8√3/2 = 4√3

OpenStudy (anonymous):

sorry that took so long :/

OpenStudy (anonymous):

it's okay! one more I promise! In a 45° - 45° - 90° triangle, the length of the hypotenuse is which of the following? a) 3 times the length of one leg b) 2 times the length of one leg c) ½ times the length of one leg d) 1 times the length of one leg

OpenStudy (anonymous):

@kittenlover731

OpenStudy (anonymous):

There are two angles of 45 which means that the triangle is isosceles. We can mark the two legs as x. We also know that according to the Pythagorean theorem, in this right triangle x² + x² = hypotenuse². SO... hypotenuse = √2x² = √2√x² = x√2 = sqrt of 2 times the length on one leg

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

yep :D

OpenStudy (anonymous):

Anything else?

OpenStudy (anonymous):

thank you so much! Well I have a question but I don't want to bother you.

OpenStudy (anonymous):

well I would try to answer it but I have a d in language arts and I have to get to work so :/

OpenStudy (anonymous):

well i'll put it in if you can answer it. & let me know if it will take to long. Find the value of x. a) 4 b) 6 c) 10 d) 16

OpenStudy (anonymous):

& ew english. I hate english. I hope there is someone here that can help you with that!

OpenStudy (anonymous):

im sorry but I really need to get to work so I cant help :( sorry @allibricker12